OpenStudy (anonymous):

27 ducks are marching to the pond. 5 lose their way, 13 return and 9 of them make it to the pond. What happens to the rest of them?

Parth (parthkohli):

5 + 13 + 9 = 18 + 9 = 27 No more ducks are left.

OpenStudy (anonymous):

yup:)